OpenStudy (anonymous):

Help me solve someone??? m=154t/(154-25t)

OpenStudy (jdoe0001):

looks "solved" already, unless you meant solve for "t"

OpenStudy (anonymous):

well i will message you the full thing then i guess

OpenStudy (jdoe0001):

tis easier if you just post in the channel, if I dunno.. then someeone else may, thus

OpenStudy (anonymous):

(1/14 + 1/11 + 1/M) * t = 1 (11/154 + 14/154 + 1/M) * t = 1 (25/154 + 1/M) * t = 1 (25/154 + 1/M) = 1/t 1/M = 1/t - 25/154 1/M = (154 - 25t) / (154t) M = 154t / (154 - 25t) is it fully solved??

OpenStudy (jdoe0001):

are you solving for "m"?

OpenStudy (anonymous):

yes

OpenStudy (jdoe0001):

ok

OpenStudy (jdoe0001):

is it fully solved?? \(\large \checkmark \)